OpenStudy (anonymous):

For any real number a > 0, with a not equal to 1, which of the following properties is not always true? A. a^x is a unique for any real number x B. a^b = a^c if and only if b = c C. For m < n, we have a ^m < a^n. D. None of the above Last one, friends. I am tallying my answers in my homework with yours. thank you so much for the help! :)

OpenStudy (hoblos):

the answer is B

OpenStudy (anonymous):

why?

OpenStudy (hoblos):

B is logically true if b=c a^b = a^b replace b=c in the right part so a^b = a^c

OpenStudy (anonymous):

but what about item A?

OpenStudy (hoblos):

for any real number x!! so if you take x=2 then x=3 is a^2 = a^3 ?! for a not equal to 1

OpenStudy (anonymous):

alright. thanks so much, hoblos. :)

OpenStudy (hoblos):

glad to help XD